(1 point) Independent random samples, each containing 90 observations, were selected from two populations. The samples from populations 1 and 2 produced 37 and 30 successes, respectively. Test H 0 :( p 1 − p 2 )=0 H0:(p1−p2)=0 against H a :( p 1 − p 2 )≠0 Ha:(p1−p2)≠0 . Use α=0.05 α=0.05 .

(a) The test statistic is

(b) The P-value is

(c) The final conclusion is

A. We can reject the null hypothesis that ( p 1 − p 2 )=0 (p1−p2)=0 and accept that ( p 1 − p 2 )≠0 (p1−p2)≠0 .

B. There is not sufficient evidence to reject the null hypothesis that ( p 1 − p 2 )=0 (p1−p2)=0 .

please show step by step and what function to use on a calculator . thank you
Answer #1

a)

p1cap = X1/N1 = 37/90 = 0.4111
p1cap = X2/N2 = 30/90 = 0.3333
pcap = (X1 + X2)/(N1 + N2) = (37+30)/(90+90) = 0.3722


Test statistic
z = (p1cap - p2cap)/sqrt(pcap * (1-pcap) * (1/N1 + 1/N2))
z = (0.4111-0.3333)/sqrt(0.3722*(1-0.3722)*(1/90 + 1/90))
z = -0.31


b)


P-value Approach
P-value = 0.7566
As P-value >= 0.05, fail to reject null hypothesis.

c)

B. There is not sufficient evidence to reject the null hypothesis that ( p 1 − p 2 )=0 (p1−p2)=0